OpenStudy (anonymous):

solve 26-13 + 17 -(-5)

OpenStudy (anonymous):

26-13+17-(-5) 26-13+17+5 13+17+5 30+5 35

OpenStudy (anonymous):

8x-3=21 8x=24 x=24/8 x=3

OpenStudy (anonymous):

yes techically to get rid of -3 you must add +3 but you have to do the same operation to both sides 8x-3+3=21+3 the yes to get the variable x by its self you must divide 8x by 8 an you must do the same operation to each side 8x/8=24/8 so yes you were right to find x you are trying to get it by itself

OpenStudy (anonymous):

Listen, I really appreciate your help...I am trying really hard to understand and I need to turn in a lot of problems on monday, so I am asking for help

OpenStudy (anonymous):

-3983-(-4123) -3983+4123 140

OpenStudy (anonymous):

thats okay im sure you can get the hang of it ... dnt forget a negative plus a negative equals a positive

OpenStudy (anonymous):

please solve -6-12-(-19)-(-8)=

OpenStudy (anonymous):

do you want to try to solve it?

OpenStudy (anonymous):

how did you get 6-12-6-(-19)+8 from -6-12-(-19)-(-8) -6-12-(-19)-(-8) -6-12+19+8 -18+19+8 1+8 =9

OpenStudy (anonymous):

solve-6,457-(-8,567)=-6,457 + 8,567=now I am stuck!

OpenStudy (anonymous):

no you simplified that right! -6457+8567= 2110

OpenStudy (anonymous):

yes when i learned this i was taught same signs add and keep different signs subtract take the sign of the larger number it will be exact so two negatives equals a negative -1+-2=-1-2=-3 positive plus a negative is which ever number is larger 9-5=4 or -19+6=-13 actually its 7222 but im sure that was a simple mistake

OpenStudy (anonymous):